OpenStudy (anonymous):

What is the equation of the line in slope–intercept form that passes through the point (0, 3) and is perpendicular to the line y = 2x – 6?

OpenStudy (anonymous):

y=-1/2 x + 3

OpenStudy (anonymous):

Nice dude. I dont understand where to go after you graph the line

OpenStudy (anonymous):

if two lines are perpendicular so we have this equation between the slope so \[\frac{m}{m \prime}=-1\]

OpenStudy (anonymous):

ooooo i c

OpenStudy (anonymous):

in this equation y=ax+b that b is y-intercept so b = 3 and we have the slope that is -1/2 that we got from the slopes equation so the tottal equation is y =-1/2 x +3

OpenStudy (anonymous):

ok cool. I understand better. Thanks alot! Can you help me witha another one?

OpenStudy (anonymous):

On the graph of the equation 6x – y = –1, what is the value of the y-intercept?

OpenStudy (anonymous):

you gotta make y alone in one side so 6x – y = –1 should be like this : y=6x+1 so y-intercept is +1

OpenStudy (anonymous):

Hell yeah dawg thanks. Do you just like doing math? Wanna do somemore?

OpenStudy (anonymous):

maybe .

OpenStudy (anonymous):

lmfaooo alright hahah What is the equation in point-slope form of the line passing through (4, 0) and (2, 6)?

OpenStudy (anonymous):

put them in this equation :\[\frac{y-y1}{y2-y1}=\frac{x-x1}{x2-x1}\]

OpenStudy (anonymous):

And then just solve?

OpenStudy (anonymous):

omg\[\frac{y-0}{6-0}=\frac{x-4}{2-4}\] \[\rightarrow -2y =6x-24 \rightarrow y=3x-12\]

OpenStudy (anonymous):

THANKS DUDE!

OpenStudy (anonymous):

Y=3(x-4)?

OpenStudy (anonymous):

no it's not necessary even it make the equation harder

OpenStudy (anonymous):

oo ok. can you reduce your answer?

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

like y=3x-12 is not an option. Which is why im thinking the answer is y=-3(x-4)

OpenStudy (anonymous):

both of 'em are true .

OpenStudy (anonymous):

oh alright thankss

OpenStudy (anonymous):

u r wllcome